OpenStudy (anonymous):

6x+2y=24 . Solve For x . So Far I Got . 6x=24+2y . Then I Got 6x=12+y . Now Im Stuck ._. Would I Divide The 6 Into The 12 ?

OpenStudy (cgreenwade2000):

Ehhhh... Check your first step again.

OpenStudy (amistre64):

you added 2y to on side, but not the other

OpenStudy (amistre64):

when they say "move it to the other side", they dont mean literally

OpenStudy (anonymous):

6x=24+2y . Now What ? & Then What Do I Do ?

OpenStudy (amistre64):

go back as step 6x+2y=24 ; what would you do to this?

OpenStudy (anonymous):

Move The 2y To The Other Side ? Leaving It Negative .

OpenStudy (amistre64):

thats not a very mathical way of stating it. we would subtract 2y from each side to keep things equal 6x + 2y = 24 - 2y -2y ------------ 6x + 0 = 24 - 2y giving us: 6x = 24 - 2y

OpenStudy (amistre64):

now, to solve for x, whats left to do?

OpenStudy (anonymous):

Divide The 6 In To Both 24 & 2y ?

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

divide each side by 6 6x 24 - 2y -- = ------ 6 6 this gives us: x = 4 -\(\large\frac13\)y

OpenStudy (anonymous):

But Isnt 6/2 3 ?

OpenStudy (amistre64):

6/3 is 2, but that has no bearing on the problem does it?

OpenStudy (amistre64):

lol, got my 2 and 3 mixed up ... but same thought

OpenStudy (anonymous):

Yeah it Does Lol .. Cause I Have To Divide 6 Into 2 .

OpenStudy (amistre64):

you have to divide 2 by 6 ....

OpenStudy (anonymous):

Oh .. Lol .

OpenStudy (amistre64):

\[\frac 26=\frac13\]

OpenStudy (anonymous):

Okaay I Get It Now ^.^ !!

OpenStudy (amistre64):

;) good luck

OpenStudy (anonymous):

Thank You Sooooooooo Much !!